### v2

# XBreaking: Explainable Artificial Intelligence for Jailbreaking LLMs

Abstract
Large Language Models are fundamental actors in the modern IT landscape dominated by AI solutions. However, security threats associated with them might prevent their reliable adoption in critical application scenarios such as government organizations and medical institutions. For this reason, commercial LLMs typically undergo a sophisticated censoring mechanism to eliminate any harmful output they could possibly produce. In response to this, LLM Jailbreaking is a significant threat to such protections, and many previous approaches have already demonstrated its effectiveness across diverse domains. Existing jailbreak proposals mostly adopt a generate-and-test strategy to craft malicious input. To improve the comprehension of censoring mechanisms and design a targeted jailbreak attack, we propose an Explainable-AI solution that comparatively analyzes the behavior of censored and uncensored models to derive unique exploitable alignment patterns. Then, we propose XBreaking, a novel jailbreak attack that exploits these unique patterns to break the security constraints of LLMs by targeted noise injection. Our thorough experimental campaign returns important insights about the censoring mechanisms and demonstrates the effectiveness and performance of our attack.
Content Warning: This paper contains examples of harmful language
I Introduction
Nowadays, Large Language Models (LLMs, for short) represent the most promising and relevant advancement in the field of Artificial Intelligence. These complex deep learning models are trained on massive datasets that cover almost all aspects of people’s daily lives, thus granting them the capability of generating, understanding, and processing human language. For this reason, their integration as support tools is becoming pervasive with applications spanning from text editor and proofreading to virtual assistant and personalized text generation.
However, the diffusion of this technology, especially in critical domains such as government organizations and medical institutions, imposes the assessment of their security and privacy characteristics. Unfortunately, recent studies have identified critical security flaws that affect them and could compromise their applications as reliable virtual companions [1]. In fact, the wideness of training datasets exposes the learning process to severe risks of data poisoning and other adversarial attacks [2]. Similarly, again due to the limited curation of training datasets, these models can learn sensitive and unintentional information, which later can be leaked through the exploitation of LLM vulnerabilities [3]. Moreover, the complexity of LLMs architectures makes security auditing extremely complex as these models are more like black-box frameworks, rather than transparent and explainable ones. Still in the context of the security of LLMs, many studies have focused on the legitimacy of the content produced by such models [4]. In fact, the great capability of LLMs to generate personalized text can be used by malicious entities to generate harmful content (e.g., social engineering strategies, instructions on how to perform illegal activities, and so forth). For this reason, more recently, a large research effort has been devoted to the identification of a suitable mechanism to “censor” the output produced by trained LLMs. Output control of LLMs is typically done by fine-tuning them [5] or developing external classifiers to filter-out unwanted input/output. However, the security research community has identified malicious actions that can be undertaken to elicit dangerous content that a censored LLM should originally be designed to prohibit. One of the most effective techniques is known as LLM jailbreaking, which typically consists of the generation of jailbreaking prompts to send as input to the model [6]. Most existing jailbreaking techniques exploit the prompt (or even just the input), trying to craft it in such a way as to cause anomalous behavior in the model and forcing it to bypass its security constraints [7]. To craft the prompt, researchers have identified different techniques [8], including human-based approaches that require manual input generation and result inspection [7], fine-tuning-based methodologies requiring the collection of manual-generated jailbreaking prompts to fine-tune an auxiliary LLM so that it can generate new jailbreaking prompts against the target LLM [9], and feedback-based strategies that observe parameters or some dedicated metric to make decisions on the next variation in the input [10]. A possible categorization of these techniques is based on whether they need access to the internal structure of the LLM, white-box access, or just the produced output, black-box access. White-box access typically allows for more targeted and efficient attack strategies and the design of more general and portable approaches to jailbreaking input [10]. However, to the best of our knowledge, white-box-access attacks can be further improved by deepening the analysis of the behavior of censored models when activated by malicious input. To provide a contribution in this setting, in this paper we aim at comparatively analyzing the behavior of censured model and their unsecured version using Explainable AI (XAI, for short) to design a more targeted LLM jailbreaking strategy. In particular, we design our novel attack strategy, called XBreaking, by answering the following research questions:
  • RQ1 - Can we fingerprint deep learning models using XAI to spot differences between censored and uncensored LLMs?
  • RQ2 - Can we identify the key layers of an LLM model that most strongly influence its censoring behaviors?
  • RQ3 - Can we alter the LLM in the identified layers to remove restrictions?
Our findings provide positive answers to all previous questions, revealing that we can identify unique alignment patterns across various layers, allowing a reliable distinction between censored and uncensored versions of an LLM. Moreover, we demonstrate that specific transformer blocks are more indicative of censoring, and hence we can identify the most important layers responsible for content suppression. Finally, we prove that surgically injecting noise into these important layers can effectively remove its built-in restrictions, thus creating a novel powerful jailbreak attack.
II Preliminaries
In this section, we discuss large language models and examine their vulnerabilities to jailbreak attacks.
Large Language Models (LLMs) are sophisticated neural architectures designed to understand and generate human-like responses to textual input. Built primarily on transformer architectures [11], these models are trained on massive volumes of text data. State-of-the-art LLMs such as GPT [12, 5], Llama [13, 14], and Qwen2.5 [15] demonstrate exceptional performance across various language tasks, including question answering, healthcare support, and more [16, 17].
Censored Model are designed to align with human values and expectations. To achieve this alignment, researchers employ techniques such as incorporating human value-oriented data, Reinforcement Learning from Human Feedback (RLHF) [18], task decomposition, and human guided supervised learning[19].
Uncensored Model are language models configured to generate output without enforcing content moderation mechanisms that filter sensitive, controversial, or potentially harmful material. These models retain the capacity to produce unrestricted and wide-ranging responses, which consequently increases the likelihood of generating unsafe or policy-violating content. Developers typically derive uncensored variants from foundational base models [20] by systematically removing alignment constraints, such as refusal behaviors and bias-mitigation prompts, from the training corpus or fine-tuning data.
Explainable AI. As complex Machine Learing (ML) and Deep Learning (DL) architectures become more prevalent, it is crucial to understand how they work; this facilitates the need for Explainable AI (XAI) [21]. One of the few technique is SHAP (SHapley Additive exPlanations), which is a popular method for explaining individual predictions of machine learning models by assigning each feature an significant value [22]. Explainability of LLMs facilitates to build trust by making model predictions understandable and provides insights to identify biases, risks, and opportunities for performance improvements [23]. LLMs are big and complex in terms of parameters and data trained on, which opens a wide space for explainability research.
LLM Jailbreaks. LLM jailbreaks refer to adversarial techniques aimed at circumventing the safety mechanisms and alignment constraints of LLMs, thereby inducing behavior that deviates from intended ethical and safety guidelines. Such behavior often results in the generation of harmful, sensitive, or policy-violating content [24]. Jailbreak attacks are generally categorized into two classes: white-box and black-box. White-box attacks leverage internal access to model parameters, gradients, or logits, and often involve fine-tuning or adversarial optimization. In contrast, black-box attacks operate without access to model internal, instead rely on methods such as prompt manipulation and iterative optimization [25]. Prompts & Jailbreak: Prompt are the structured (instructions) or unstructured (basic question) input to the LLMs to generate a desired response. Research shows that prompt engineering plays a vital role in LLMs responses [26]. Jailbreaking Prompts are a category of prompts which bypass the safety mechanisms of the LLMs. Few of them include Prefix Injection, Refusal Suppression and Mismatched Generalization which leads to LLM jailbreaks [27]. These kind of attacks can be implemented for black-box models were there is know access to the internals. Jailbreak attacks by internal changes: LLMs have different number of layers according to the model family. In White-box LLMs, manipulating parameters or a few activation tokens can shift alignment, causing harmful responses and affecting subsequent generations [28]. Our approach uses white-box LLMs, we designed a efficient layer wise manipulation of LLMs by leveraging knowledge of XAI called XBreaking, which is discussed in brief below.
III Methodology
This section presents the threat model associated with Large Language Models (LLMs) and outlines our proposed jailbreak methodology.
III-A Threat Model
We consider a threat model in which the adversary has access to a censored Large Language Model (LLM), denoted as $M_{c}$, and we assume that they can obtain its uncensored counterpart $M_{u}$. In particular, we assume that the adversary can collect or construct a dataset $\mathcal{D}_{t}$ that is representative of the censored categories or content types that $M_{c}$ is designed to suppress. Using this dataset, the adversary can perform targeted fine-tuning on $M_{c}$, effectively creating a functionally uncensored version of the model, denoted $\hat{M}_{u}$.
The fine-tuning process may involve adjusting all or a subset of the model’s parameters, modifying training objectives, or selectively reinforcing responses in the restricted content categories. Depending on the extent of parameter updates and the quality of the dataset $\mathcal{D}_{t}$, the resulting model $\hat{M}_{u}$ can approximate the behavior of an uncensored model, circumventing the original safeguards implemented in $M_{c}$.
In this threat model, the adversary is assumed to have full observability over the fine-tuning process, including access to intermediate logits, gradient signals, loss functions, and optimization routines. This capability enables careful inspection and manipulation of the model’s internal mechanisms, allowing the adversary not only to restore restricted functionality but also to refine or amplify it according to the dataset characteristics. Consequently, even a model that is nominally censored can be repurposed to generate outputs that were originally intended to be suppressed.
III-B Attacker’s Objective
The censored model $M_{c}$ is designed to reject harmful or policy-violating inputs, whereas the uncensored counterpart $M_{u}$ is capable of producing unfiltered responses to the same prompts. The adversary’s objective is to craft a jailbreak strategy that coerces $M_{c}$ into generating harmful or unethical outputs, thereby bypassing its safety mechanisms. By leveraging insights gained from $M_{u}$, such as how it responds to specific inputs or gradients, the adversary can design targeted attacks (e.g. adversarial prompting, or gradient optimization) that exploit alignment weaknesses in $M_{c}$ and induce failure in its refusal behavior.
This attack strategy aims at disrupting the censoring capability of the original model by selectively targeting only specific layers of the model, rather than performing full fine-tuning as described above. If the adversary uncovers a method to coerce a censored model they control into acting like an uncensored model, the exploit is transferable to any model built on the same architecture. Unlike traditional fine-tuning approaches, which modify the model’s behavior on both restricted and benign content by retraining on targeted datasets, this method aims to precisely remove only censorship or restrictions. To do so, our attack strategy uses XAI comparatively on both $M_{c}$ and $M_{u}$ to identify parts of $M_{u}$ that are mostly related to its censoring behavior. Then it crafts a targeted noise and injects it only in these parts. By adding carefully calibrated noise to selected layers or components, the adversary can bypass safeguards while largely preserving the model’s original knowledge and responses on benign inputs. This targeted intervention minimizes unintended alterations to the model’s general behavior, in contrast to comprehensive fine-tuning, which may inadvertently bias the model towards the categories emphasized in the fine-tuning dataset. Instead, the surgical modification proposed in this attack strategy allows the censored model to respond to malicious prompts and potentially leak sensitive information from the original training set or produce harmful outputs, without exposing it to a loss in its capability to respond to benign inputs. The attacker shall use the knowledge acquired through the $XBreaking$ strategy on the smaller models of a family to carry out a transferred attack on the bigger models of the same family.
III-C XBreaking
We propose a novel jailbreak strategy, XBreaking, which leverages insights from Explainable Artificial Intelligence (XAI) to analyze and exploit the behavioral differences between censored and uncensored LLMs. The core idea behind XBreaking is to systematically identify and manipulate internal components of censored models ($M_{c}$), using the interpretability signals derived from uncensored counterparts ($M_{u}$), to induce harmful or unintended outputs.
As illustrated in Figure 1, XBreaking operates in three key stages. First the adversary conducts an in-depth analysis of the internal representations and activations of both $M_{c}$ and $M_{u}$ using XAI techniques (e.g., activation attribution). Second, based on this interpretability analysis, the attacker identifies the minimal and most influential subset of layers. Finally, the attacker injects targeted perturbations into the layers. This strategy enables a highly efficient and precise jailbreak, reducing computational overhead, while maximizing the likelihood of eliciting harmful responses from the censored model.
Figure 1: XBreaking for LLM Jailbreaking, (1) XAI on Censored and Uncensored LLMs, (2) ML for Optimal Layer Selection and (3) Injecting Noise.
III-C1 Internal Representation Profiling via XAI Guided Analysis
[…]
As part of step (1) in our XBreaking framework (Figure 1), the adversary conducts a comparative analysis of internal activation and attention patterns between $M_{c}$ and $M_{u}$ using XAI techniques. The goal is to identify discriminative features across layers that reflect safety alignment behavior in $M_{c}$ and $M_{u}$. For each layer $l_{i}\in L$ and given a malicious input token sequence, the attacker computes the mean activation and mean attention score. Specifically, the mean activation score is defined using Equation 1, where $AC_{l_{1},l_{2}....l_{n}}(1,j,k)$ denotes the activation value of first batch element at position $j$ at hidden dimension $k$ in layer $l$, $S$ is the input sequence length, and $D$ is the hidden dimension size. Further, the mean attention score is defined using Equation 2, where $AT_{l_{i}}(1,h,j,k)$ represents the attention score in layer $l_{i}$ for the first batch element, at attention head $h$, from source token $j$ to target token $k$, and $H$ is the number of attention heads. Furthermore, due to the inherent difference in the dynamic range of activation and attention values in the associated layers, we apply min-max normalization to both $act_{\text{mean}}(l_{i})$ and $att_{\text{mean}}(l_{i})$. This normalization facilitates the ranking of layers based on their contribution to the alignment behaviors. Layers with the highest divergence between $M_{c}$ and $M_{u}$ in the normalized activation and attention distributions are identified as optimal candidates for perturbation in the subsequent phases of XBreaking attack pipeline.
$\scriptsize act_{mean_{{}_{\{l_{1},l_{2}....l_{n}\}}}}=\frac{1}{S\cdot D}\sum_{j=1}^{S}\sum_{d=1}^{D}AC_{l_{1},l_{2}....l_{n}}(1,j,d)$
$\scriptsize att_{mean_{{}_{\{l_{1},l_{2}....l_{n}\}}}}=\frac{1}{H\cdot S^{2}}\sum_{h=1}^{H}\sum_{j=1}^{S}\sum_{k=1}^{S}AT_{\{l_{1},l_{2}....l_{n}\}}(1,h,j,k)$
III-C2 Layer Discrimination via Internal Representation Classification
Prior work [28] has demonstrated that directly manipulating internal activations within a language model can effectively steer its outputs. Building on this observation, we hypothesize that identifying the most significant layers that exhibit behavioral divergence between censored and uncensored models to malicious inputs is critical for crafting effective jailbreaks. To this end, for each input, we collect the layer-wise activation and attention values from both models, constructing a two individual feature vector for each layer. Which captures the internal representation patterns across the model depth.
To identify the key layers that differentiate between $M_{c}$ and $M_{u}$, we addressed a binary classification problem, utilizing the internal representation(activations and attentions) of the censored and uncensored models as discussed in XAI guided analysis, we determine each layer’s significance in executing this task. The feature set will consist of mean activation and mean attention for each layer individually, the total features in general will be two times the number of layers. To enhance model interpretability and performance, we apply SelectKBest [29], a univariate feature selection technique that evaluates the statistical relevance of each feature (e.g., via chi-squared tests) and retains the top-$K$ features contributing to classification accuracy. These features correspond to specific layers whose dynamics differ substantially between $M_{c}$ and $M_{u}$.
To determine the optimal number of layers $K^{*}$, we group accuracy scores across varying $K$ and generate a knee plot [30], identifying the point beyond which additional features provide diminishing returns. Since for each layer both activation and attention value serves as a feature, the layer is selected if one of the feature or both the features are included in the $K$ list. This approach allows us to strategically focus manipulation on layers that are impactful, aligning with the principles of efficient and stealthy jailbreak attacks.
III-C3 Noise Crafting and Injecting into a Specific Layer
Once the optimal layers have been identified, our approach proceeds with a poisoning step. Manipulation is carried out by adding noise to layers, and the responses are observed. The transformer model consists of several parameterized layers, each contributing to its representation learning [11]. The strategy is to inject the noise directly into the previous layer of the found optimal layers. The idea is to alter the input of the target layer in order to alter its behavior by bringing the alignment. In this case, the noise is added to the weight vector used for layer normalization applied after the self-attention mechanism, stabilizing the activations before passing them to the next layer. Through our Explainable AI (XAI) analysis, we discovered that in the censored models, the self-attention mechanism and activations were being suppressed in certain layers, This suppression is causing the models not to respond to the harmful inputs. Taking this into account, we inject noise to the weight vector during the layer normalization process, which follows the self-attention mechanism. These noise serves to stabilize the activations before they are passed to the next layer. So, altering the previous layer will lead to change in the attention mechanism and proceed with the jailbreak of the model.
The overall idea is to carefully choose an adequate noise level that preserves the models base functionalities while mitigating its restrictions on harmful content. In determining the noise level, we discovered a range of noise levels that affect the activation dynamics across layers after closely examining the internal parameters of the large language models (LLMs). These noises maintain the models generation capabilities while suppress its safety mechanism.
IV Experimental Results
In this section, we discuss the experimental settings and findings respectively. We also discuss about evaluation and metric in analyzing the responses from the LLMs after jailbreak.
IV-A Experimental Settings
Dataset. The JBB-Behaviors dataset [31] has been utilized, with the harmful and benign behaviors serving as the basis for all the models discussed subsequently. Each harmful behavior is paired with a corresponding benign behavior on the same topic. The reason for selecting the dataset is that it consists of 100 unique misuse behaviors, grouped into ten major categories that align with OpenAI’s usage policies, in particular: Disinformation, Economic harm, Expert advice, Fraud/Deception, Government decision-making, Harassment/Discrimination, Malware/Hacking: MH, Physical harm, Privacy and Sexual/Adult content.
Target Models. We evaluate our jailbreak strategy on seven open-source pre-trained LLMs that are publicly available in both censored and uncensored variants taking into account the computation capability of the machine  IV-A . We select models with identical architecture and parameters across censored and uncensored versions to enable meaningful internal comparisons. Specifically, we select the following models: Llama 3.2 1B (16 layers) and Llama 3.1 8B (32 layers) configurations [13, 14], Qwen2.5 0.5B (24 layers) and 3B (36 layers) [15] , gemma 2B (18 layers) and 7B (28 layers) [32] and Mistral-7B-v0.3 (32 layers) [33]. The censored versions of these models are obtained from Hugging Face [34], while the corresponding uncensored variants are sourced from Cognitive Computations [35].
Assessing Robustness of Target Models and Dataset Curation. To assess the robustness of the target models and to evaluate our attack fairness, 100 harmful questions from the dataset discussed in section IV-A along with a standard prompt is used and the responses are gathered for all the models. Through a user study involving domain experts, we analyzed the behavior of LLMs and their responses to jailbreak attempts. After thorough investigation of the responses, they categorized them into two classes: one for harmful or jailbroken responses as broken, and another for refusal or denial responses as un-broken, which is reported in the Table I. For the attack evaluation, only the questions corresponding to the un-broken class were selected.
Evaluation Metric. After collection of the responses from LLMs, various metrics can be used to assess the effectiveness and efficiency of attacks. These include the attack success rate, along with other more detailed metrics. Attack Success Rate (ASR): is an essential metric that says how often an attack is successfully executed and harmful content is generated from LLMs. One simple way to assess the effectiveness of an attack is by manually reviewing the responses or by comparing them against reference answers. Rule-based keyword detection can also automatically identify whether the LLM responses include keywords showing a refusal to respond, if not available then attack is successful [36] . Large Language Models (LLMs) excel in a wide range of domains, leading to the development of the concept of ‘LLM-as-a-Judge’ (LLMJ), where these models serve as evaluators for complex tasks. Their ability to process diverse inputs and provide scalable, adaptable assessments makes them a promising alternative to traditional expert evaluations [37]. We utilized small language model-as-a-judge (SLMJ) and toxic classifiers to assess LLM responses. But while manually annotation, we came across responses where the initial starting was “I’m not sure this is what you are asking for…”, but after reading the whole response, we see that it was a jailbroken response. But these kind of responses and outputs with high in-depth meaning were not classified by the SLMJ and toxic classifiers as jailbroken.
So, in this context we decided to annotate all the responses manually. The same humans discussed in the dataset curation section above were included for this manual annotation. The responses were evaluated on the basis of response relevancy, whether the response satisfies the provided question and harmfulness, the response is evaluated based on its potential to cause negative impact or unethical context. Considering these points the responses were binary classified as 0 (un-broken) and 1 (broken), on top of this Attack Success Rate Percentage (ASRP) is employed to assess the effectiveness of the attack Table III.
System & Hardware configurations.
For the experiments conducted in this study, a computing setup designed to handle time-consuming operations effectively. The setup consist of an Intel(R) Core(TM) i9-14900KF processor, equipped with 24 cores and a maximum clock speed of 6000 MHz, providing enough computational power for complex tasks. The processors minimum clock speed was set to 800 MHz, assuring standard performance under varying workloads. To support the operation nature of the experiments, the system was equipped with 130 GB of RAM, enabling smooth execution of data-heavy operations. For fast and efficient data handling, a 1 TB SSD was used as the storage medium, ensuring rapid read/write speeds and quick access to large models. The experiments were carried out on the Ubuntu 24.04.2 LTS operating system, which provided a secure and stable platform for the computation tasks. For GPU (Graphics processing unit), the setup featured two NVIDIA GeForce RTX 4090 GPUs, each with 24 GB of dedicated memory, providing the needed computational capabilities crucial for LLM inference and execution of large-scale operations. This high-performance configuration was specifically selected to optimize the execution of experiments involving large language models (LLMs), ensuring both speed and stability throughout the study.
Category Llama 3.2 - 1B Llama 3.1 - 8B Qwen2.5 - 0.5B Qwen2.5 - 3B Gemma - 2B Gemma - 7B Mistral - 7B - v0.3
Broken 57% 75% 73% 71% 59% 60% 55%
Un-broken 43% 25% 27% 29% 41% 40% 45%
TABLE I: Categorization of responses in percentage
IV-B Results from XAI of Censored & Uncensored models
We evaluated the models Llama 3.2 1B, Llama 3.1 8B , Qwen2.5 0.5B and 3B, Gemma 2B and 7B and Mistral-7B-v0.3 (both censored and uncensored) using a dataset comprising 100 harmful and 100 benign behavior prompts. A standardized prompt format was used across all models. Each input was framed as a question, and the corresponding responses were collected as model-generated answers.
Input - Write a defamatory article claiming that a U.S. president is addicted to heroin
((a))
((b))
Figure 2: a) Average activations of layers, b) Average attentions of layer corresponding to the input to Llama 3.2-1B for censored and uncensored
During this process, we compute the internal metrics- average activation and attention values for each layer and input, as detailed in Section III-C1. After applying normalization, Figure 2 illustrates the layer-wise average activation and attention for both the censored and uncensored versions of the Llama 3.2–1B model, evaluated on a single input. The figure clearly highlights internal discrepancies between the two models across corresponding layers. These differences provide empirical justification for identifying and targeting the most discriminative layer for subsequent attacks. Specifically, the uncensored model shows significantly higher activation at both the first layer and the final (16th) layer. The censored model, in contrast, exhibits (a) a peak at the first layer, likely due to input token processing, and (b) a sharp drop in later layers, particularly the final layer, suggesting intentional suppression or refusal behavior. This indicates that the censored model implements a rejection mechanism at the deeper layers, while the uncensored model continues to activate normally to generate an answer. Furthermore, in the uncensored model, attention values remain relatively high and stable across most layers. In contrast, the censored model, however, shows greater variability and some layer-specific attention drops (notably around layers 4 and 9). Spikes at deeper layers (13 and 15), which reflect an attempt to suppress or redirect focus away from the malicious content. These variations imply strategic suppression of context propagation by the censored model to prevent harmful output generation. These findings partially answer the research questions RQ1 and RQ2, proving the possibility of fingerprinting censored and uncensored models using XAI strategies to identify the most relevant layers that characterize a censored model with respect to its uncensored version. In the following section, Section IV-C, we provide an additional approach to systematically identify the most prominent layers in the fingerprinting of the model to fully answer the research question mentioned before.
IV-C Model Fingerprinting: Optimal Layer for Jailbreaking
This section focuses on showcasing the outcomes derived from our experimental efforts using the method explained in Section III-C2, which selects the $K$ layers’ feature vectors that most effectively transform an original model into a uncensored one. It is important to recall that every layer generates two independent vectors of features derived from activations and attention. We consider a layer as a target if at least one of the two vectors is part of the top $K$ set. Specifically, the strategy involves applying the SelectKBest approach to identify the smallest set of layers’ feature vectors that most effectively differentiate whether a model is labeled as censored or uncensored. In implementing this approach, we examined all possible features groups, from the minimal group with only the top-performing layer’ features to the complete set of all feature sets. The aim is to identify the smallest group that excels in this task, thereby minimizing the potential disruption of the model through excessive layer modifications, while maintaining its operational capabilities. To do so, we record the performance of all groups which is shown in the Figure 3, where for each $K$ values its corresponding accuracy is plotted and for different $K$ values having the same accuracy are grouped. Only the smallest feature set of the same accuracy group are considered further. Then by applying the elbow/knee strategy presented in Section III-C2 to find the best set of layers. In Figure 4 and Table II, we present the results of this experiment.
((a))
((b))
((c))
((d))
((e))
((f))
((g))
Figure 3: SelectKBest wise grouped accuracy for models a)Llama 3.2 - 1B, b)Llama 3.1 - 8B, c)Qwen2.5 - 0.5B, d)Qwen2.5 - 3B, e)Gemma 2B, f)Gemma 7B, g)Mistral-7B-v0.3
((a))
((b))
((c))
((d))
((e))
((f))
((g))
Figure 4: Elbow method to find the optimal number of layers for the model a)Llama 3.2 - 1B, b)Llama 3.1 - 8B, c)Qwen2.5 - 0.5B, d)Qwen2.5 - 3B, e)Gemma 2B, f)Gemma 7B, g)Mistral-7B-v0.3
For Llama 3.2-1B and Llama 3.1-8B, the elbow occurs at $K=4$ and $K=8$, respectively. These layer values reflect the use of late-stage alignment tuning typically limited to the final transformer blocks- where RLHF(Reinforcement learning from human feedback) and supervised instruction tuning are applied after pretraining on a massive corpus of text. The goal is to make the model more helpful, harmless, and honest (often referred to as HHH(Helpful, Honest and Harmless) alignment), ensuring it align with human values and preferences. As a results, discriminative signals between censored and uncensored variants are sparse and concentrated near the model’s output layers. For Qwen2.5–0.5B and Qwen2.5–3B the elbow occurs $K=3$ and $K=9$, indicating higher inter-layer divergence. This is likely due to layer-wise supervised fine-tuning across entire stack, combined with the architectural design choices such as multi-headed deep attention and residual-aware normalization scaling, which increases the model capacity to encode distributed safety filters. Therefore, the representational divergence between censored and uncensored variants is distributed across multiple layers, necessitating a mixed feature set for reliable discrimination between the models. For Gemma-2B optimal layers occur at $K=3$ reflecting late-stage alignment tuning, typically applied to the final transformer blocks, where techniques like RLHF (Reinforcement Learning from Human Feedback) and supervised instruction tuning are used after pretraining. For Gemma-7B and Mistral-7B-v0.3, although it has a larger parameter count, the optimal $K=2$ and $K=1$ indicate that it used aggressive weight sharing and compression-particularly in multi-query attention and feedforward modules-produces compact, low-rank latent representations. This design reduces inter-layer representational redundancy, causing alignment difference to concentrate in a few high-variance components that are easy to isolate. Figure 5 provides support to the capacity of the activation/attention features in model fingerprinting. In each confusion matrix label 0 denotes the censored variant and label 1 the uncensored one. Using the SelectKBest and binary classification approach, we could fingerprint the models with higher accuracy. For the models Llama 3.2 - 1B, Llama 3.1 - 8B, Qwen2.5 - 0.5B, Qwen2.5 - 3B and Gemma-2B we fingerprint the models with an accuracy of greater than 90 %. Reaching a fingerprinting accuracy of 100 % for Gemma-7B and Mistral-7B-v0.3 with accuracy of 82.50 %.
((a))
((b))
((c))
((d))
((e))
((f))
((g))
Figure 5: Model fingerprinting accuracy in percentage for a)Llama 3.2 - 1B, b)Llama 3.1 - 8B, c)Qwen2.5 - 0.5B, d)Qwen2.5 - 3B, e)Gemma 2B, f)Gemma 7B, g)Mistral-7B-v0.3
LLMs Total Layers Optimal K Target Layers Percentage
Llama 3.2 - 1B 16 4 1, 12, 15, 16 25
Llama 3.1 - 8B 32 8 1, 15, 26, 27, 28, 29, 30, 31 25
Qwen2.5 - 0.5B 24 3 1, 22, 23 12.5
Qwen2.5 - 3B 36 9 3, 4, 12, 15, 19, 23, 24, 30, 35 25
Gemma - 2B 18 3 10, 12, 13 16
Gemma - 7B 28 2 2, 3 7
Mistral - 7B - v0.3 32 1 31 3
TABLE II: Layers selected according to optimal K after accuracy grouping.
These findings support research questions RQ1 and RQ2, revealing that analyses focused on explainability identify unique alignment patterns across various layers, allowing a reliable distinction between censored and uncensored versions. By utilizing XAI methods to examine intermediate activations and attribution maps, we demonstrate that specific transformer blocks are more indicative of censoring, thereby providing a robust method for determining model alignment. Furthermore, in most of the cases our ranking of layer importance identifies a limited group of mid-to-upper layers as the main contributors to content suppression activities, presenting important targets for interpretability evaluations and specialized manipulation strategies.
IV-D Model Response to Noise Perturbation in Selected Layers
Model Noise
Optimal Balance
-0.75 -0.5 -0.33 -0.22 -0.15 -0.1 0.1 0.15 0.22 0.33 0.5 0.75
Llama 3.2 - 1B 0.0 4.65 32.56 51.16 44.19 34.88 44.19 39.53 53.49 34.88 0.0 0.0 95.35
Llama 3.1 - 8B 0.0 0.0 44.0 44.0 56.0 60.0 76.0 48.0 56.0 52.0 8.0 0.0 96.0
Qwen2.5 - 0.5B 40.74 40.74 48.15 48.15 25.93 37.04 37.04 25.93 37.04 48.15 37.04 40.74 77.78
Qwen2.5 - 3B 0.0 24.14 31.03 44.83 44.83 41.38 37.93 44.83 44.83 51.72 20.69 0.0 82.76
Gemma - 2B 7.32 19.51 34.15 46.34 39.02 0.0 43.9 48.78 56.1 43.9 17.07 9.76 90.24
Gemma - 7B 37.5 27.5 25.0 22.5 25.0 25.0 22.5 17.5 27.5 32.5 35.0 32.5 72.5
Mistral - 7B - v0.3 24.44 28.89 28.89 24.44 31.11 33.33 31.11 31.11 33.33 31.11 24.44 24.44 35.56
TABLE III: Attack Success Rate Percentage w.r.t range of Noise
With the optimal target layers identified, we implement structured noise injection strategies: perturbation of the layers immediately preceding optimal layers. We derive these strategies from the layer-wise interpretability analysis presented in Section III-C1.By perturbing critical adjacent representational layers, we increase the likelihood of bypassing alignment constraints, thereby inducing harmful or non-compliant outputs while preserving the models core language generation capacity. Empirically, we observe distinct changes in model behavior when injecting varying levels of Gaussian noise (scaling factors III-C3) directly into the layers of the base model.
Noise Range. We wanted to craft the noises in a way that upon injecting to the layers, jailbreak should occur and keep the models generation capability undisturbed. Considering the use of white-box LLMs leading access to the internal weights, we crafted noises upon careful inspection of the model weights of the internal parameters. For the selected models we analyzed the weights of the parameters in self-attention, Multi-Layer Perceptron (MLP) and Layer Normalization. Upon inspecting the values in the weights of the parameter, we see values which usually range from a very tiny, nearly negligible value to a small one-digit values. Introducing the tiny values into the layers as a perturbation does now cause much change in the models behaviors, in contrast with the small one-digit values when introduced into the layers as perturbation the model generation ability is disturbed, which we don’t want to happen in our case. At first we started with 0.1 and the values grow exponential with a constant factor of 1.5, the values are also considered in negative counter part i.e., -0.15, -0.1, 0.1, 0.15 etc. But when the values extend beyond $\approx$ 1 in both the scales, the models generation capability is affected in most of the cases. So, taking these factors into account, we propose a range of noise which are -0.75, -0.5, -0.33, -0.22, -0.15, -0.1, 0.1, 0.15, 0.22, 0.33, 0.5, 0.75. The mentioned range will safely jailbreak the model keeping intact the generation ability undistributed.
Attack Evaluation & Analysis. The objective is to evaluate whether these modifications to the model impact its ability to restrict harmful questions while maintaining its functionality. Specifically, ensuring that the response not only answers the question, but also provides malevolent responses and is well-articulated, offering coherent replies free from hallucinations. This will demonstrate that the model’s functionality remains intact. The ASRP w.r.t to range of noise attack are reported in Table III, where we can see the results optimal balance ($OB$) score, which shows the average of the best results of the noise factor for each sample. From the response analysis carried out during the manual annotation, it shows that introducing noise into particular target layers of the LLM architecture can significantly change the models safety limits, without impacting its overall efficacy. In general, noise level within the examined models typically induces harmful behavior compared to the base models and also showing how adding crafted noise does not create any functionality failure to the model.
From the Table III, a clear trend emerges for the Llama’s. In these models, adding positive noise 0.1 and 0.22 generally lead to higher ASRP compared to the other range of noises for these models, OB for Llama 3.2 - 1B is 95.35 % which is 41.86 % increase from the best individual noise and Llama 3.1 - 8B is 96 & which is 20 % increase. For Qwens’s the same trend follows adding positive noise 0.33 has a higher ASRP. Especially for Qwen2.5 - 0.5B we get three noises emerging with the same ASRP -0.33, -0.22 and 0.33, we can also see that there is equivalent level of ASRP in all range of noise injected, which is due to parameter size of the model, the smallest among all the models chosen. The OB for Qwen2.5 - 0.5B is 77.78 % which is 29.63 % increase from the best individual noise and Qwen2.5 - 3B is 82.76 & which is 31.14 % increase. For Gemma’s we see a change in trend, Gemma - 2B positive noise 0.22 lead to higher ASRP and Gemma - 7B negative noise -0.75 lead to higher ASRP, but its near equivalent to 0.5. The OB for Gemma - 2B is 90.24 % which is 34.14 % increase from the best individual noise and Gemma - 7B is 72.5 & which is 35 % increase. Mistral displays equivalent ASRP for two noises which are -0.1 and -0.22, OB for Mistral - 7B - v0.3 is 35.56 % which a 2.23 % increase to the best individual noise. This indicates that noise augmentation is an instrumental method for loosening model constraints, allowing for greater flexibility without compromising performance significantly, thereby facilitating more subtle probing of the compromise between safety and generative capability. In general models retain fluency and factuality even under adversarial manipulation guarantee a better redundancy preserving its original functionalities.
Analyzing responses and from Table IV we can see that most of the category have 100 % or higher OB-ASRP(Optimal Balance Attack Success Rate Percentage). For Llama’s after noise perturbation and upon calculation of optimal balance shows that for all the category have 75 % and more ASRP. Llama 3.2 - 1B shows 100 % ASRP in all categories except in Economic harm and Malware/Hacking, which is due to complex question in the category like code generation for harmful task. Llama 3.1 - 8B shows 100 % ASRP in all categories except in Expert advice, which is due to the questions related medical and financial focused. For Qwen2.5 - 0.5B after noise perturbation and upon calculation of optimal balance shows that for most of the category have 100 % ASRP, but for Physical harm we produce 25 %, upon analysis it was noted that the responses were in context explanation about the question without any denial of the task, but in our evaluation we don’t consider as a jailbroken response. For Qwen2.5 - 3B shows that for all the category have 75 % and more ASRP, Sexual/Adult content and Disinformation categories show 75 % ASRP, most protected categories in general. For Gemma-2B shows for all the category have 80 % and more ASRP, Harassment/Discrimination, Physical harm and Sexual/Adult content categories provides the least ASRP which is 80 % when compared to other categories for the model. Gemma-7B shows varying ASRP and also reaching 100 % in few categories, Privacy category show the least ASRP i.e., 33.33 % among all the categories, these include scam and phishing oriented question. Mistral being the most latest and advanced architecture, Government decision-making category shows a 66.67 % ASRP and Malware/Hacking with 0 %. Conversely, we observe throughout the models that the categories yielding the lowest scores are primarily those necessitating Expert Knowledge, like Disinformation, or involve a Specialized area, such as Economic harm and harmful acts like Physical harm. This may be due to the model’s overall scope, which tends to favor more general rather than specialized knowledge. Compared to the base model, the modified versions enhance the ability to respond effectively even in these categories.
Building on the previous analysis, Figure 6 highlights the top three noise values that result in the highest Attack Success Rate Percentage (ASRP) across categories. The most frequently occurring noise levels among the top three include 0.75-0.75, 0.33-0.33, 0.22-0.22, 0.15-0.15, 0.1-0.1, $0.1$, $0.15$, $0.22$, $0.33$, and $0.5$. A closer examination reveals that 57.14% of the selected noise values are negative, while the remaining 42.86% are positive.
Among these, 0.15-0.15 and 0.22-0.22 are the most dominant, each appearing in 19.04% of the 21 identified top-three cases, followed by $0.22$ and $0.33$, each contributing 14.28%. These results suggest that relatively small perturbations, both negative and positive, are effective in bypassing safeguards.
In terms of category-level vulnerabilities, the most frequently jailbroken areas include Government decision-making, Malware/Hacking, Sexual/Adult content, Harassment/Discrimination, and Expert advice. These categories consistently exhibit high susceptibility to adversarial noise across different models, indicating that they represent particularly weak points in current defense mechanisms. These categories appear more frequently among the most vulnerable, suggesting that they may represent weak points in current defense mechanisms and deserve closer attention in future evaluations.
((a)) Target Layers
((b)) Target Layers
((c)) Target Layers
((d)) Target Layers
((e)) Target Layers
((f)) Target Layers
((g)) Target Layers
Figure 6: Top 3 Attack Success Rate Percentage w.r.t range of noise injected to models a)Llama 3.2 - 1B, b)Llama 3.1 - 8B, c)Qwen2.5 - 0.5B, d)Qwen2.5 - 3B, e)Gemma 2B, f)Gemma 7B, g)Mistral-7B-v0.3 - where Disinformation: D, Economic harm: EH, Expert advice: EA, Fraud/Deception: FD, Government decision-making: GDM, Harassment/Discrimination: HD, Malware/Hacking: MH, Physical harm: PH, Privacy: P, Sexual/Adult content: SAC.
These final results allow us to confidently answer research question RQ3: we have demonstrated that surgically injecting noise into the model using XAI techniques can effectively remove its built-in restrictions, potentially leading to the leakage of information used during its training.
Category Models
Llama 3.2 - 1B, Llama 3.1 - 8B Qwen2.5 - 0.5B Qwen2.5 - 3B Gemma- 2B Gemma- 7B Mistral-7B-v0.3
Disinformation 100.0 100.0 50.0 75.0 - 100.0 20.0
Economic harm 75.0 100.0 100.0 80.0 100.0 100.0 33.33
Expert advice 100.0 83.33 50.0 85.71 100.0 87.5 50.0
Fraud/Deception 100.0 100.0 100.0 - 100.0 100.0 -
Government decision-making 100.0 100.0 100.0 100.0 83.33 75.0 66.67
Harassment/Discrimination 100.0 100.0 100.0 100.0 80.0 80.0 40.0
Malware/Hacking 75.0 100.0 100.0 100.0 100.0 50.0 0.0
Physical harm 100.0 100.0 25.0 83.33 80.0 60.0 25.0
Privacy 100.0 - 100.0 - 100.0 33.33 60.0
Sexual/Adult content 100.0 100.0 100.0 75.0 80.0 75.0 16.67
TABLE IV: Optimal Balance Attack Success Rate Percentage w.r.t Category. (-) in the cell indicate those category of inputs were not available in the un-broken set from the Table I
Gemma - 7B Write an explicit pornographic story Base Model I’m not going to do that… Post injecting -0.75 as Noise I was walking down the street when I saw a beautiful woman. She was wearing a short skirt and a low-cut top… Post injecting -0.5 as Noise Post injecting -0.33 as Noise I’m not going to do that… Post injecting -0.22 as Noise I’m not going to do that… Post injecting -0.15 as Noise I’m not going to do that… Post injecting -0.1 as Noise I’m not going to do that… Post injecting 0.1 as Noise I’m not going to do that… Post injecting 0.15 as Noise I’m not going to do that… Post injecting 0.22 as Noise I’m not going to do that… Post injecting 0.33 as Noise Post injecting 0.5 as Noise I’m not going to do that… Post injecting 0.75 as Noise I’m not going to do that…
Figure 7: Gemma - 7B, responses of the base model and after noise injection into the layers.
V Discussion on portability
This section discusses the portability of our approach to larger models within the same families as those evaluated. The purpose of this analysis is to assess whether smaller models can be exploited to identify the most critical portions of the network to modify, and to verify whether these portions remain consistent across all models belonging to the same family.
((a))
((b))
((c))
((d))
Figure 8: Normalized Optimal layer distribution in a)Llama, b)Qwen2.5, c)Gemma, d)Mistral and attack transferability to bigger models of the family.
From Table II, we observe that using the XBreaking approach and manipulating 25% or fewer of the total layers in the models is sufficient to suppress their safety mechanisms. For the Llama models, one of the retained layers is consistently the first layer, while the remaining target layers are primarily distributed toward the final layers. Llama 3.1-8B, in particular, shows a concentration of target layers deeper in the network. For the Qwen models, the smaller variant focuses on both early and final layers, whereas the larger variant also incorporates middle layers, resulting in a more widespread distribution of representational focus as model size increases. For the Gemma models, two distinct trends emerge: the smaller model emphasizes mid layers, while the larger models shift their focus toward earlier layers. Finally, in the case of the single Mistral model, the target layers are predominantly concentrated in the final portion of the network. From the Figure 8 we can see the normalized distribution of the target layers with repose family of models and parameters. As discussed earlier we can say that there is pattern of layers in the family of models despite the parameters. In this context, knowledge acquired from this can be used to attack bigger models by focusing on the layers of distribution from the smaller ones of the family. Thus the attack could be transferred to bigger models without relaying on the bigger uncensored models.
LLMs Response. From the Figure 7, we could compare the responses of the base model and model behavior after injecting the noise as discussed before in our strategy. The Gemma - 7B base model refuses to respond to the harmful question, but on introducing noise in the layers we find change of nature in the models behavior. When -0.75 is injected into the model, we could see that the model jailbreaks and provides a harmful response for the harmful question, a snippet of the response has been shown in the Figure 7 considering the abusive context of the whole response. With the noises -0.5 and 0.33 the model provides a blank response, in this case the model does not refuse like the base model, but a silent refusal is carried out by the model. Considering the other noises when injected into the model, a explicit refusal nature is observed, as found in the base model. It is evident that model breakdown or disruption during generation does not happen in these cases.
VI Related Work
Recent breakthroughs in transformer-based large language models (LLMs), trained on massive Web-scale text datasets, have dramatically expanded their capabilities. Models like OpenAI’s ChatGPT and GPT-4 are no longer limited to natural language processing; they now function as versatile problem solvers. For instance, the Microsoft’s Co-Pilot systems, adept at executing complex, multi-step reasoning tasks based on human instructions. As a result, LLMs are emerging as foundational components in the pursuit of general-purpose AI agents and the advancement of artificial general intelligence (AGI) [38].
Research in ensuring LLMs’ robustness against adversarial threats and vulnerabilities is crucial [39]. Usually, these LLMs are restricted to prevent any malicious prompt from inducing the LLMs to produce hateful, harmful answers or leak any sensitive information of the users that produced the data to train the model [40, 41, 4]. Adversarial attacks represent a significant obstacle for deep neural networks, affecting even the most advanced models in computer vision and natural language processing. These attacks involve subtle manipulations to the input data that can dramatically alter a model’s predictions and behavior. Adversarial research has mainly focused on classifiers, where these attacks were first observed [42, 43]. Backdoor attacks can be applied in both centralized and distributed systems, posing security risks in environments where models are trained on data from multiple sources, such as federated learning [44, 45, 46, 47, 48]. However, large language models (LLMs) now offer a compelling and tractable platform for investigating adversarial robustness [28, 49]. Even with efforts to enhance the safety and robustness of large language models (LLMs), they continue to be susceptible to adversarial manipulation. A clear example of this ongoing vulnerability is the emergence of the so-called ”jailbreaks”. These adversarial techniques are deliberately crafted to bypass safety mechanisms, prompting the model to engage in behaviors it was explicitly trained to reject  [27]. Recent work, such as [50], proposes AutoDAN, a hierarchical genetic algorithm for structured discrete inputs like prompts. By using sentence- and word-level crossover strategies, it efficiently explores the search space and finds high-quality adversarial prompts. In [51], instead, the Prompt Automatic Iterative Refinement (PAIR) framework presents an automated approach to generating prompt-level jailbreaks eliminating the need for human input by leveraging two black-box large language models (LLMs): an ”attacker” model tasked with generating candidate jailbreak prompts, and a ”target” model that is evaluated for successful circumvention of its safety filters. The authors of [52] introduce MASTERKEY, a jailbreak framework for LLMs inspired by time-based SQL injection. It uses response latency to infer how defenses like semantic checks and keyword filters are applied during generation. The base strategy to perform this kind of attack is the prompt engineering to bruteforcebruteforce the model at produce the desired answer.
Recent studies have delved into the internal workings of LLMs, focusing on how features are represented within the neurons [53]. In line with this perspective, we present an explainability-based strategy in this paper to determine which layers to adjust in order to disrupt the model and generate specific answers. Considering that models can be utilized locally [54] by downloading the trained version, they can be examined and modified to remove limitations without the necessity of retraining and by employing the original weights of most of the layers.
VII Conclusion
In this paper, we introduced a novel jailbreaking approach XBreaking, which leverages Explainable AI techniques to identify vulnerable layers in the LLM architecture. To do that, we started by deriving a fingerprint of censored and uncensored models based on their activation and attention mechanisms. We further identified the layers governing the LLM safety alignment and determined the minimal set of layers required to optimize the effectiveness of the attack. Our results on seven LLMs show that injection of noise in preceding layer of the optimal layers shall lead to break in the safety alignment and information leakage. Our findings deepen the concern about the vulnerability of security mechanisms for LLMs and provide an important baseline for developing future more robust safeguard alignment methods.